OpenStudy (anonymous):

MEDAL!!! The leg of a right triangle is 3 units and the hypotenuse is 11 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (jhannybean):



OpenStudy (jhannybean):

For special right triangles, we can apply The Pythagorean Theorem, which states, \((\text{hypotenuse})^2 = (\text{leg 1})^2 + (\text{leg 2})^2\)

OpenStudy (alexandervonhumboldt2):

by pyth theorem x^2+3^2=11^2

OpenStudy (jhannybean):

Or it can rather be written as \(c^2 = a^2+b^2\) where c is the hypotneuse, and a and b are the legs. Now if we are given the length of the hypotenuse and a leg of a triangle, we can rearrange our equation to solve for the leg of the triangle

OpenStudy (alexandervonhumboldt2):

x^2=121-9

OpenStudy (alexandervonhumboldt2):

x=sqrt(112)

OpenStudy (jhannybean):

Therefore, \(11^2 = 3^2 +x^2 \implies x = \sqrt{11^2 -3^2}\)

OpenStudy (jhannybean):

Now just solve for x and you will find the length of the missing leg :)

OpenStudy (anonymous):

so its 

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

ok